I am having trouble with square roots. Can you help me with these problems, please?
1. Multiply -- sqrt 3/5 x sqrt 6/4 =
2. Simplify -- sqrt 27/p^2 =
3. Divide -- sqrt 9x^5y divided by sqrt 12x^2y^6 =
4. Divide -- 2 divided by sqrt 3 plus sqrt 6 =
Thank you.

Answer by stanbon(75887)

1. Multiply -- sqrt (3/5) x sqrt (6/4) = sqrt(18/20) = (3/2)sqrt(2/5)
-------------------------------------------------------------------
2. Simplify -- sqrt (27/p^2) = (3/p)sqrt(3)
-------------------------------------------------------------------
3. Divide -- sqrt 9x^5y divided by sqrt 12x^2y^6
= sqrt[(3/4)x^3/y^5]
= (x/y^2)sqrt(3x/y)
-------------------------------------------------------------------
4. Divide -- 2 divided by sqrt 3 plus sqrt 6
= 2/(sqrt(3)+sqrt(6))
---
Multiply numerator and denominator by(sqrt(3)-sqrt(6)) to get:
---
= (2(sqrt(3)-sqrt(6))/(3-6)
---
= (-3/2)(sqrt(3)-sqrt(6))
=============================
Cheers,
Stan H.
=============================